OpenStudy (anonymous):

When attempting to do his homework, Mark incorrectly uses the addition property of equality below. Which statement would correct Mark’s error? 1) 8 + x = 2 2) 8 – 8 + x = 2 3) x = 2

OpenStudy (john_es):

\[8-8+x=2-8\Rightarrow x=-6\] Mark must substract the same number to both members of the equation.